OpenStudy (anonymous):

Which statement shows the distributive property? A. 13 · (k + 10) = 13 · k + 13 · 10 B. 13 · (k + 10) = 13 · k + 10 C. 13 · (k + 10) = 13 · k · 10 D. 13 · (k + 10) = (k + 10) · 13

OpenStudy (akashdeepdeb):

A. 13 · (k + 10) = 13 · k + 13 · 10 Because, Distributive property goes like this! a(b+c) = ab + ac. :D Understood? :)

OpenStudy (anonymous):

ya thanks.

OpenStudy (akashdeepdeb):

:)